OpenStudy (anonymous):

Which of the following is a third root of the given complex number? 27(cos(pi/5)+isin(pi/5)) Check all that apply. A. 3(cos(8pi/15)+isin(8pi/15)) B. 3(cos(11pi/15)+isin(11pi/15)) C. 27(cos(11pi/15)+isin(11pi/15)) D. 3(cos(pi/15)+isin(pi/15))

OpenStudy (ikram002p):

27(cos(pi/5)+isin(pi/5))^1/3 hint :use this identity \(( r (\cos \theta+i \sin \theta ))^n= r^n (\cos n\theta+i \sin n\theta ) \)

OpenStudy (ikram002p):

27^1/3(cos(pi/5)+isin(pi/5))^1/3 = 27^1/3 (cos(pi/15)+isin(pi/15) ) =3 (cos(pi/15)+isin(pi/15) )

OpenStudy (ikram002p):

got it ? let me know if u still have some doubts